# Local consistency and axioms of functional dependence

### v1

@@ Abstract
Local consistency arises in diverse areas, including Bayesian statistics, relational databases, and quantum foundations. Likewise, the notion of functional dependence arises in all of these areas. We adopt a general approach to study logical inference in a setting that enables both global inconsistency and local consistency. Our approach builds upon pairwise consistent families of $K$-relations, i.e, relations with tuples annotated with elements of some positive commutative monoid. The framework covers, e.g., families of probability distributions arising from quantum experiments and their possibilistic counterparts. As a first step, we investigate the entailment problem for functional dependencies (FDs) in this setting. Notably, the transitivity rule for FDs is no longer sound, but can be replaced by two novel axiom schemes. We provide a complete axiomatisation and a PTIME algorithm for the entailment problem of unary FDs. In addition, we explore when contextual families over the Booleans have realisations as contextual families over various monoids.
@@ 1 Introduction
Local consistency is an important concept in various areas such as Bayesian statistics, relational databases and quantum foundations. In broad terms, local consistency refers to a family of partial structures (such as marginal distributions or projections) that agree on their overlapping parts. Local consistency is usually a desirable property to enforce or at least a minimal requirement or observed phenomenon. It is often contrasted with global consistency, which demands the existence of a single underlying structure whose projections yield the given family. In practice, local consistency sometimes serves as a proxy for global consistency, especially when the latter is infeasible to check or maintain explicitly.
In Bayesian statistics, local consistency is central to the representation of global probability distributions via factorisation. For example, Bayesian networks encode a global distribution as a family of marginal and conditional distributions that must be consistent with each other. Such factorisations are very useful as they facilitate computationally feasibly probabilistic and approximate reasoning without the need to directly access or store the global distribution.
@@
In relational databases, local consistency arises when data is spread across multiple tables with overlapping attribute sets. The motivation here is not computational speed-up—storing everything in a single table would eliminate costly joins in query evaluation—but data integrity, as decomposing data helps to prevent anomalies, such as update anomalies. Notably, consistency in relational databases is typically enforced via integrity constraints that are local in nature. E.g., data about a person may be distributed across multiple tables, which we expect to be globally consistent; yet this consistency is usually maintained via primary keys and foreign keys that act on individual tables or pairs of tables only.
@@
In quantum theory, local consistency is a principle that places an upper bound on the outcomes appearing in physical experiments: any experimental setup must produce a family of probability distributions that satisfies local consistency. However, not all locally consistent families are compatible with quantum predictions.
This intricate relationship between local and global consistency naturally gives rise to many interesting questions. In domains where local consistency is used as a proxy for global consistency, it is natural to ask: under what conditions does local consistency entail global consistency? If local consistency is a necessary but not sufficient condition for observable phenomena, one may query about the reasons that account for this gap.
For instance, quantum theory offers remarkably accurate predictions on experimental outcomes, yet it may not provide a satisfactory explanation—in terms of fundamental principles or laws of nature —as to why certain locally consistent distributions are not physically realisable. Several efforts have been made to explain this discrepancy in terms of simple fundamental principles. A notable example is the principle of information causality (?), which uses the language of information theory to constrain admissible distributions.
@@
In the context of databases and probability distributions, in turn, much effort has gone into characterising when local consistency guarantees global consistency. The so-called local-to-global consistency refers to the situation where the underlying structure (such as a database schema or collection of variable sets) guarantees that every locally consistent family (of relations or distributions) can be extended to a globally consistent one. The characterisations of this property are typically formulated in the language of graph theory. For relational databases it is known that a database schema satisfies the local-to-global consistency property if and only if it forms an acyclic hypergraph (?). Similar characterisations have been also shown in the context of distributions and multisets (??). More recently, these results have been generalised in the setting of $K$-relations, which are relations whose tuples are annotated with elements from some positive and commutative monoid (?). In this case, the assumption of acyclicity of the schema is not always sufficient but a more involved characterisation is required.
Our Contributions. In this article, we turn the tables and treat local consistency not as a proxy for global consistency, but as a foundational concept in its own right. Analogously to (?), we adopt the useful abstraction of $K$-relations that allows us to treat families of relations and distributions uniformly. We also redirect our focus from domains such as information theory and graph theory to logical consequence, and investigate how it is impacted by the shift from global to local consistency. As a test case, we study the logical entailment between functional dependencies (FDs). Our approach bears similarity to the logic of local inference of ? (?) and to quantum team logic of (?) which are geared for applications in quantum theory, where local consistency does not generally entail the existence of a global model (cf. Bell’s theorem).
@@
After defining our theoretical framework, we investigate when a given family of locally consistent relations $\mathcal{R}$ can be enriched to a family of locally consistent $K$-relations (with the same support). Here, we give a complete characterisation for the case, where $K$ is cancellative and $\mathcal{R}$ is a contextual family over what we refer to as a chordless-cycle context set. Moreover, we establish a sound and complete axiomatisation for the entailment of unary FDs in the setting of locally consistent $K$-relations. In particular, we observe that the usual transitivity rule of FDs becomes unsound in this context. Interestingly, the transitivity rule disperses into an infinite collection of weaker rules that we name chain and cycle rules. In the standard unirelational context these rules are all derivable from the transitivity rule, and one instance of a chain rule derives the transitivity rule. Finally, we establish that the derivability question of our axiomatisation can be decided in polynomial time. As an instantiation of our result, we obtain insight on the entailment of FDs in settings that lack global consistency, such as in quantum foundations.
@@ 2 Preliminaries
For a natural number $n$, we write $[n]:=\{0,1,\dots,n-1\}$. We fix a countably infinite set Var of variables. Given a set $A$, an assignment (of $A$) is a function $s$ that maps a finite set $D\subseteq\textrm{Var}$ of variables to some values (in $A$). We call $D$ the domain of $s$ and write $\mathrm{Dom}(s)$. The set of all assignments $s\colon D\to A$ is denoted $\textrm{As}(D,A)$.
A monoid is an algebraic structure $K=(K,+,0)$, where $+$ is associative and 00 is the identity element of $+$. $K$ is positive if $a+b=0$ entails $a=0=b$, for all $a,b\in K$, and (left) cancellative if $a+b=a+c$ entails $b=c$. We associate each monoid $K$ with its natural order $\leq$, defined by $a\leq b\iff\exists c:a+c=b$. The natural order of a monoid is reflexive and transitive, meaning that it is a preorder. In this paper, $K$ will always denote a positive commutative non-trivial monoid (the trivial monoid contains only 00).
@@
where $s\!\restriction\!D^{\prime}$ is the restriction of the assignment $s$ to $D^{\prime}$ and $\sum$ is the aggregate sum of the monoid.
In addition to the general case, we consider the Boolean monoid $\mathbb{B}=(\{0,1\},\lor,0)$ and the monoids of non-negative reals $\mathbb{R}_{\geq 0}=([0,\infty),+,0)$ and natural numbers $\mathbb{N}=(\mathbb{N},+,0)$ with their usual addition.
3 $K$-relations and Contextual Families
In quantum information theory, a context is a set of attributes that can be measured together. More formally, a context $C$ is a set of variables and context set $\mathcal{C}$ is a downward closed set of contexts (i.e, $C\in\mathcal{C}$ and $C^{\prime}\subseteq C$ implies $C^{\prime}\in\mathcal{C}$). We often represent a context set $\mathcal{C}$ in terms of its maximal elements; for instance, we may write $\mathcal{C}=\{\{x,y\},\{y,z\}\}$ to denote $\mathcal{C}=\{\{x,y\},\{y,z\},\{x\},\{y\},\{z\},\emptyset\}$. To remove clutter, we often write a context $\{x_{1},\dots,x_{n}\}$ as $x_{1}\dots x_{n}$.
@@ 3 $K$-relations and Contextual Families
Let $\mathcal{C}$ be a context set. A contextual $K$-family over $\mathcal{C}$ is a locally consistent set of $K$-relations $\mathcal{R}$ containing exactly one $K$-relation $R$ with domain $C$, for each $C\in\mathcal{C}$. The $K$-relation of $\mathcal{R}$ with domain $C$ is referred to as $\mathcal{R}_{C}$. The set $\mathcal{C}$ is the context set of $\mathcal{R}$, denoted $\textrm{Conts}(\mathcal{R})$.
The support $\mathrm{Supp}(\mathcal{R})$ of $\mathcal{R}$ is defined as $\{\mathrm{Supp}(\mathcal{R}_{C})\mid C\in\mathcal{C}\}$. A contextual $\mathbb{B}$-family is simply called a contextual family, and identified with its support.
@@
If $\mathcal{R}$ is a contextual family and $\alpha\in K$, we set $\alpha\mathcal{R}$ as $\{\alpha\mathcal{R}_{C}\mid C\in\mathcal{C}\}$. In contrast to the above, if $\mathcal{R}$ is a contextual family, $\alpha\mathcal{R}$ is not necessarily a contextual $K$-family (it may violate local consistency; cf. Ex. 5). Let $\mathcal{R}$ and $\mathcal{S}$ be contextual $K$-families over a shared context set $\mathcal{C}$. We set $\mathcal{R}+\mathcal{S}$ as $\{\mathcal{R}_{C}+\mathcal{S}_{C}\mid C\in\mathcal{C}\}$. Clearly, local consistency is preserved under this operation.
@@
If are $\mathcal{R}$ and $\mathcal{S}$ are contextual $K$-families over a shared context set $\mathcal{C}$, then $\mathcal{R}+\mathcal{S}$ is a contextual $K$-family.
A contextual family $\mathcal{R}$ is sometimes identified with the set of its assignments $\bigcup{\mathcal{R}}=\bigcup_{C\in\mathcal{C}}\mathcal{R}_{C}$. Similarly, a contextual $K$-family $\mathcal{S}$ can be viewed as a single function #:Supp($\#:\bigcup{\mathrm{Supp}(\mathcal{S})}\to K_{\neq 0}$, where $K_{\neq 0}=\{a\in K\mid a\neq 0\}$. A contextual family is called $K$-realisable if it is the support of some contextual $K$-family.

### v2

# Locally Consistent K-relations: Entailment and Axioms of Functional Dependence

@@ Abstract
Local consistency arises in diverse areas, including Bayesian statistics, relational databases, and quantum foundations, and so does the notion of functional dependence. We adopt a general approach to study logical inference in a setting that enables both global inconsistency and local consistency. Our approach builds upon pairwise consistent families of $K$-relations, i.e, relations with tuples annotated with elements of some positive commutative monoid. The framework covers, e.g., families of probability distributions arising from quantum experiments and their possibilistic counterparts. As a first step, we investigate the entailment problem for functional dependencies (FDs) in this setting. Notably, the transitivity rule for FDs is no longer sound, but can be replaced by two novel axiom schemas. We provide a complete axiomatisation for, and establish NL-completeness of, the entailment problem of unary FDs, and demonstrate that even this restricted case exhibits context-dependent subtleties. In addition, we explore when contextual families over the Booleans have realisations as contextual families over various monoids.
@@ 1 Introduction
Local consistency is an important concept in various areas such as Bayesian statistics, relational databases and quantum foundations. In broad terms, local consistency refers to a family of partial structures (such as marginal distributions or projections) that agree on their overlapping parts. Local consistency is usually a desirable property to enforce, or at least a minimal requirement for an observed phenomenon. It is often contrasted with global consistency, which demands the existence of a single underlying structure whose projections yield the given family. In practice, local consistency sometimes serves as a proxy for global consistency, especially when the latter is infeasible to check or explicitly maintain.
@@
In relational databases, local consistency becomes an issue when data is spread across multiple tables with overlapping attribute sets. The motivation for using multiple tables is data integrity; decomposing data helps prevent anomalies, e.g., related to updates. Notably, consistency in relational databases is typically enforced via integrity constraints that are local in nature. For example, data about a person may be distributed across multiple tables, which we expect to be globally consistent; yet this consistency is usually maintained via primary and foreign keys that act on individual tables or pairs of tables only.
@@
In quantum theory, local consistency is a principle that places an upper bound on the outcomes appearing in physical experiments: any experimental setup must produce a family of probability distributions that satisfies local consistency (a.k.a. no-signalling). However, not all locally consistent families are compatible with quantum predictions (see, e.g,  (?)) and, vice versa, there are quantum realisable measurement scenarios that induce locally consistent probability distributions that are not globally consistent (?; ?). This intricate tension between local and global consistency (contextuality) has been extensively studied in quantum information theory (see, e.g., (?; ?; ?)) and has been recognised as a fundamental resource in quantum computation, providing a key explanation for why quantum systems can outperform classical ones (?).
In Bayesian statistics, local consistency is central for the representation of global probability distributions via factorisation. For example, Bayesian networks encode a global distribution as a family of marginal and conditional distributions that must be consistent with each other. Such factorisations are very useful as they facilitate computationally feasible probabilistic and approximate reasoning without the need to directly access or store the global distribution (see, e.g, (?)).
@@
The intricate relationship between local and global consistency raises natural questions: under what conditions does local consistency entail global consistency. The so-called local-to-global consistency refers to the situation where the underlying structure (such as a database schema or a collection of variable sets) guarantees that every locally consistent family (of relations or distributions) can be extended to a globally consistent one. The characterisations of this property are typically formulated in the language of graph theory. For relational databases it is known that a database schema satisfies the local-to-global consistency property if and only if it forms an acyclic hypergraph (?). Similar characterisations have been also shown in the context of distributions and multisets (?; ?). More recently, these results have been generalised in the setting of $K$-relations, which are relations whose tuples are annotated with elements from some positive commutative monoid (?). In this case, the assumption of acyclicity of the schema is not always sufficient but a more involved characterisation is required.
A distinct but related area to study is what kind of inferences remain valid when only local consistency is guaranteed. In this setting, classical inference principles that rely on global consistency may fail. A particularly simple but instructive example is the failure of transitivity of functional dependence: even if $A$ functionally determines $B$ and $B$ functionally determines $C$ in two locally consistent views, it does not necessarily follow that $A$ functionally determines $C$ globally.
Example 1.
Patient Diagnosis
P1 Cancer
P2 ALS
Diagnosis Treatment
Cancer Chemotherapy
ALS Palliative
Patient Treatment
P1 Chemotherapy
P1 Palliative
P2 Palliative
Figure 1: Locally consistent family of relations for patients, diagnoses, and treatments.
Figure 1 illustrates three distinct semantic views to hospital data. In the top-left table, each patient is assigned exactly one diagnosis, captured by the functional dependency $\texttt{Patient}\rightarrow\texttt{Diagnosis}$, and in the bottom-left table, each diagnosis is associated with a preferred treatment according to clinical guidelines, yielding $\texttt{Diagnosis}\rightarrow\texttt{Treatment}$. The right-hand side table reflects actual clinical practise where doctors have freedom to deviate from standard guidelines, for instance by performing different treatments in sequence when earlier interventions fail. As a consequence, this table does not satisfy $\texttt{Patient}\rightarrow\texttt{Treatment}$, in violation of transitive closure. The three tables are locally consistent in the sense that any two tables agree on patients, treatments, and diagnoses whenever they record them. However, the tables are not globally consistent, since there exists no single relation whose projections coincide with all three tables simultaneously.
More generally, treatments of patients may be irreducible to any finite set of rules, even when additional variables are introduced, as there are always contextual factors and exceptions that escape formal representation. Nevertheless, we may still wish to retain certain rules as abstract guidelines in an idealized context. In our example, $\texttt{Diagnosis}\rightarrow\texttt{Treatment}$ captures a general rule in a context where individual patients are absent by design. The example thus models a situation where real-life complexity forces us to drop the assumption of global representability (analogously to the lack of global classical probability distribution in quantum physics).
The tension between local and global consistency plays a central role in Ontology-Based Data Access (OBDA), where reasoning is often performed over locally consistent views induced by ontologies and mappings, even when global consistency of the underlying data is absent or undecidable to check. Prior work (?; ?) has highlighted the practical and theoretical relevance of local reasoning principles in this setting, showing that meaningful query answering and inference can be achieved under local consistency assumptions. Other prior frameworks that develop the theory of logical inference in connection to local consistency include the logic of local inference of (?) and quantum team logic by ? (?), both of which are geared for applications in quantum theory.
Our Contributions. In this article, we turn the tables and treat local consistency not as a proxy for global consistency, but as a foundational concept in its own right. Analogously to (?), we adopt the useful abstraction of $K$-relations, where $K$ is a monoid, as they uniformly generalise Boolean relations, bags, and probability distributions, while preserving the algebraic structure (addition) needed for marginalising and reasoning about consistency. This algebraic abstraction also highlights prominent characteristics of relations, bags, and distributions, clarifies when and why they diverge, and offers a framework for generalisation to other contexts. We also redirect our focus from domains such as information theory and graph theory to logical consequence, and investigate how it is impacted by the shift from global to local consistency. As a test case, we study the logical entailment (i.e., the implication problem) between functional dependencies (FDs) over locally consistent families of $K$-relations. It is worth noting that the assumption of local consistency makes it possible to assign a truth value to an FD over a family of relations with shared attributes. It is important to note that while satisfaction of FDs on any given family of locally consistent $K$-relations depends only on the support of that family, validity of entailment can depend on the monoid, for not all monoids yield the same class of supports of locally consistent $K$-relations.
@@
After defining our theoretical framework, we establish a sound and complete axiomatisation for the entailment of unary FDs with binary CDs in the setting of locally consistent $K$-relations. A context dependency (CD) is a kind of a dummy FD of the form $x\rightarrow x$ that only declares that variables in $x$ are locally consistent. Further, we observe that the usual transitivity rule of FDs becomes unsound. Interestingly, the transitivity rule disperses into an infinite collection of weaker rules that we name chain and cycle rules. In a globally consistent setting, these rules are all derivable from the transitivity rule, and one instance of a chain rule derives the transitivity rule. We then establish that the derivability problem of our axiomatisation for unary FDs is NL-complete in conjunction with binary CDs, and remains in polynomial time with arbitrary CDs. See Table 1 for a summary.
FDs CDs Axioms Complexity Reference
[…]
unary $\dagger$ R+T NL easy observation
unary binary R+CL NL Thms. 11 & 14
unary binary single R+CL NL proof of Thm. 14
[…]
unary ternary single CH (s) NL Prop. 23
Table 1: Overview of our axiomatisation and complexity results on entailment of inclusion dependencies. R, A, and T refer to the standard Armstrong’s axioms of reflexivity, augmentation, and transitivity (see page 3.2), and CL and CH refer to the cycle and chain rules (see pages 11 and 22). Axiomatisations marked with (s) are sound, others are sound and complete. Single means that derivations can use only assumptions and a single application of a proof rule. Complexity results refer to completeness, if not otherwise stated. $\dagger$: non-contextual setting. $\ddagger$:(?; ?).
Finally, we investigate when a given family of locally consistent relations $\mathcal{R}$ can be enriched to a family of locally consistent $K$-relations (with the same support). ? (?) showed, via examples and using different terminology, that there are families of locally consistent relations that cannot be expanded to families of locally consistent probability distributions with the same support. We extend these examples to a complete characterisation of this extendability property in the case where $K$ is cancellative and $\mathcal{R}$ is a contextual family over what we refer to as a hole context sequence.
Lastly, we show that the set of supports of contextual $\mathbb{R}_{\geq 0}$-families and contextual $\mathbb{N}$-families coincide. Hence, an axiomatisation for FDs is sound (resp. complete) w.r.t. contextual $\mathbb{R}_{\geq 0}$-families iff it is that for contextual $\mathbb{N}$-families.
@@ 2 Preliminaries
For a natural number $n$, we write $n]:=\{0,1,\dots,n-1\}$. We fix a countably infinite set Var of variables. Given a set $A$, an assignment (of $A$) is a function $s$ that maps a finite set $D\subseteq\textrm{Var}$ of variables to some values (in $A$). We call $D$ the domain of $s$ and write $\mathrm{Dom}(s)$. The set of all assignments $s\colon D\to A$ is denoted $\textrm{As}(D,A)$.
A monoid is an algebraic structure $K=(K,+,0)$, where ++ is associative and 0 is the identity element of ++. $K$ is positive if $a+b=0$ entails $a=0=b$, for all $a,b\in K$, and (left) cancellative if $a+b=a+c$ entails $b=c$. We associate each monoid $K$ with its natural order $\leq$, defined by $a\leq b\iff\exists c:a+c=b$. The natural order of a monoid is reflexive and transitive, meaning that it is a preorder. In this paper, $K$ will always denote a positive commutative non-trivial monoid (the trivial monoid contains only 0). Commutativity and positivity enforce that the marginalisation of a K-relation is well-defined and that the operations of taking a support and a marginalisation commute (these operations are defined next).
@@
where $s^{\prime}\!\restriction\!D^{\prime}$ is the restriction of the assignment $s^{\prime}$ to $D^{\prime}$ and $\sum$ is the monoid’s aggregate sum. In addition to the general case, we consider the Boolean monoid $\mathbb{B}=(\{0,1\},\lor,0)$ and the monoids of non-negative reals $\mathbb{R}_{\geq 0}=([0,\infty),+,0)$ and natural numbers $\mathbb{N}=(\mathbb{N},+,0)$ with their usual addition.
In quantum information theory, a context is a set of attributes that can be measured together. More formally, a context $C$ is a set of variables, and a context set $\mathcal{C}$ is a downward closed set of contexts (i.e, $C\in\mathcal{C}$ and $C^{\prime}\subseteq C$ implies $C^{\prime}\in\mathcal{C}$). We often represent a context set $\mathcal{C}$ in terms of its maximal elements; for instance, we may write $\mathcal{C}=\{\{x,y\},\{y,z\}\}$ to denote $\mathcal{C}=\{\{x,y\},\{y,z\},\{x\},\{y\},\{z\},\emptyset\}$. To avoid clutter, we often write a context $\{x_{1},\dots,x_{n}\}$ as $x_{1}\dots x_{n}$. We call $\mathcal{C}$ a uni-context, if $\bigcup\mathcal{C}\in\mathcal{C}$.
@@
A contextual $K$-family over a context set $\mathcal{C}$ is a locally consistent set of $K$-relations $\mathcal{R}$ containing exactly one $K$-relation $R$ with domain $C$, for each $C\in\mathcal{C}$. The $K$-relation of $\mathcal{R}$ with domain $C$ is referred to as $\mathcal{R}_{C}$. The set $\mathcal{C}$ is the context set of $\mathcal{R}_{\mathcal{C}}$.
The support $\mathrm{Supp}(\mathcal{R})$ of $\mathcal{R}$ is defined as $\{\mathrm{Supp}(\mathcal{R}_{C})\mid C\in\mathcal{C}\}$. A contextual $\mathbb{B}$-family is simply called a contextual family, and identified with its support. Since $K$ is assumed to be positive, we obtain the following proposition.
@@
If $\mathcal{R}$ is a contextual family and $\alpha\in K$, we set $\alpha\mathcal{R}$ as $\{\alpha\mathcal{R}_{C}\mid C\in\mathcal{C}\}$. In contrast to the above, if $\mathcal{R}$ is a contextual family, $\alpha\mathcal{R}$ is not necessarily a contextual $K$-family (it may violate local consistency; cf. Ex. 15). Let $\mathcal{R}$ and $\mathcal{S}$ be contextual $K$-families over a shared context set $\mathcal{C}$. We set $\mathcal{R}+\mathcal{S}$ as $\{\mathcal{R}_{C}+\mathcal{S}_{C}\mid C\in\mathcal{C}\}$. Clearly, local consistency is preserved under this operation.
@@
If $\mathcal{R}$ and $\mathcal{S}$ are contextual $K$-families over a shared context set $\mathcal{C}$, then $\mathcal{R}+\mathcal{S}$ is a contextual $K$-family.
A contextual family $\mathcal{R}$ is sometimes identified with the set of its assignments $\bigcup{\mathcal{R}}=\bigcup_{C\in\mathcal{C}}\mathcal{R}_{C}$. Similarly, a contextual $K$-family $\mathcal{S}$ can be viewed as a single function $\bigcup{\mathrm{Supp}(\mathcal{S})}\to K_{\neq 0}$, where $K_{\neq 0}=\{a\in K\mid a\neq 0\}$. A contextual family is called $K$-realisable if it is the support of some contextual $K$-family.
@@
The contextual family in Fig. 1 is locally but not globally consistent (as stated in Example 1). In particular, each pair of tables agree on the projection to their common part. To see why global consistency does not hold, suppose toward a contradiction that there exists a single relation that projects to the three tables in the figure. Let $s$ be an assignment of this relation mapping $(\texttt{Patient},\texttt{Treatment})$ to (P1, Palliative). Then $s$ must map $(\texttt{Diagnosis},\texttt{Treatment})$ to (ALS, Palliative), as can be seen from the bottom-left table. Howe, the top-left table entails $s$ maps $(\texttt{Patient},\texttt{Diagnosis})$ to (P2, ALS), which means that $s$ maps Patient to two different values, a contradiction.
3 Entailment of Dependence
Next, we focus on entailment of functional dependencies on contextual $K$-families. We first summarise the well-known Armstrong’s axioms for functional dependence in the classical setting (?). We then develop the theory in the contextual setting. In particular, we give a sound and complete axiomatisation for unary dependencies and establish that the related derivability problem is NL-complete.
3.1 Classical Setting
A functional dependency (FD) is a statement of the form $x\rightarrow y$, where $x,y\subseteq D$ are finite sets of variables. An FD $x\rightarrow y$ is satisfied in a relation $X$ (written $X\models x\rightarrow y$) if for all $s,s^{\prime}\in X$: $\bigwedge_{z\in x}s(z)=s^{\prime}(z)$ entails $\bigwedge_{z\in y}s(z)=s^{\prime}(z)$. For a set $\Sigma$ of FDs, we write $\Sigma\models x\rightarrow y$ if for all relations $X$: if $X\models\phi$ for all $\phi\in\Sigma$, then $X\models x\rightarrow y$.
In (?), it is established that the entailment $\Sigma\models x\rightarrow y$ is completely axiomatised by the following set of three axioms, termed Armstrong’s axioms: